OpenStudy (anonymous):

If the volume of a pyramid is 60 cubic inches and the area of the base is 36 square inches, what is the height of the pyramid?

OpenStudy (anonymous):

5in

OpenStudy (anonymous):

or h= 5in

OpenStudy (anonymous):

They're right.

OpenStudy (anonymous):

\[V= \frac{ lwh }{ 3 }\]

OpenStudy (anonymous):

\[\frac{ 6 \times6 \times 5 }{ 3 }\]

OpenStudy (anonymous):

\[6 \times 6 \times 5=180\]\[180\div3=60\]

OpenStudy (anonymous):

So yes, the answer is 5